What is the approximate diameter of a wheel that made 752 revolutions when it rolled 1 kilometer?
a) 24.2 cm b) 32.0 cm c) 42.3 cm d) 84.6 cm e) 113 cm

Answer by richwmiller(9135)

1km=1000meters
1000 meters=1000*100 cm
1000*100/752=100,000/752=the circumference of the wheel in cm
circumference=2*pi*r=d*pi
where d=diameter and r =radius and pi=3.14